Hoodmemes:

The seventh grade enrollment at Main Middle School reduced from 615 students to 492 students. By What percentage was the number of students reduced?

Convert:

615-492=123 percent change= (decrease/original number)*100 percent change= (123/615= 0.2)*100 percent change= 20% decrease

Hoodmemes:

ok,thank you

Convert:

Happy to help! :)